OpenStudy (anonymous):

Which is an equation of the line passing through points at (1, 6) and (-2, 1)?

OpenStudy (anonymous):

make it a vector: (1,6) - (-2,1) = (3,5) (x,y) = (1,6) +k(3,5) x = 1+3k (=) k = (x-1)/3 y = 6+5k (=) y = (x-1)/3 equation line: y = (1/3)x -1/3